# SplatMAP: Online Dense Monocular SLAM with 3D Gaussian Splatting

### v2

@@ Abstract.
Achieving high-fidelity 3D reconstruction from monocular video remains challenging due to the inherent limitations of traditional methods like Structure-from-Motion (SfM) and monocular SLAM in accurately capturing scene details. While differentiable rendering techniques such as Neural Radiance Fields (NeRF) address some of these challenges, their high computational costs make them unsuitable for real-time applications. Additionally, existing 3D Gaussian Splatting (3DGS) methods often focus on photometric consistency, neglecting geometric accuracy and failing to exploit SLAM’s dynamic depth and pose updates for scene refinement. We propose a framework integrating dense SLAM with 3DGS for real-time, high-fidelity dense reconstruction. Our approach introduces SLAM-Informed Adaptive Densification, which dynamically updates and densifies the Gaussian model by leveraging dense point clouds from SLAM. Additionally, we incorporate Geometry-Guided Optimization, which combines edge-aware geometric constraints and photometric consistency to jointly optimize appearance and geometry of the 3DGS scene representation, enabling detailed and accurate SLAM mapping reconstruction. Experiments on the Replica and TUM-RGBD datasets demonstrate the effectiveness of our approach, achieving state-of-the-art results among monocular systems. Specifically, our method achieves a PSNR of 36.864, SSIM of 0.985, and LPIPS of 0.040 on Replica, representing improvements of 10.7%, 6.4%, and 49.4%, respectively, over the previous SOTA. On TUM-RGBD, our method outperforms the closest baseline by 10.2%, 6.6%, and 34.7% in the same metrics. These results highlight the potential of our framework in bridging the gap between photometric and geometric dense 3D scene representations, paving the way for practical and efficient monocular dense reconstruction. A demonstration of the results can be found in the accompanying video: https://youtu.be/Pr_kyWQQkGo.
@@
ccs: Computing methodologies Reconstruction
@@ 1. Introduction
Figure 1. Visual comparison of reconstructed scenes. Our method (SplatMap) achieves superior rendering fidelity and geometric accuracy, closely matching the ground truth, while outperforming GIORIE-SLAM (Zhang et al., 2024) and Photo-SLAM (Huang et al., 2024) in fine structural details.
@@
In order to achieve real-time 3D reconstruction with high-fidelity rendering, researchers have explored the potential of integrating SLAM with differentiable rendering (Zhu et al., 2022, 2024; Yang et al., 2022; Keetha et al., 2024; Johari et al., 2023; Sandström et al., 2023; Yan et al., 2024a; Matsuki et al., 2024; Zhang et al., 2024; Huang et al., 2024). However, monocular SLAM systems, particularly during early mapping stages, are prone to producing inaccurate point clouds due to limited observations, shallow triangulation baselines, and weak pose constraints. These inaccuracies propagate into downstream mapping tasks, leading to artifacts such as ”ghosting” on walls and furniture, especially when observed from untrained viewpoints. Traditional 3DGS densification pipelines typically rely on RGB losses to correct these errors, introducing additional computational overhead as they compensate for the initial inaccuracies.
@@ 2.1. Simultaneous Localization and Mapping (SLAM)
In contrast, dense SLAM methods aim to reconstruct detailed 3D representations of the environment, often leveraging depth maps or pixel-level photometric consistency for optimization. ElasticFusion (Whelan et al., 2015) circumvented traditional pose graph optimization by directly refining dense RGB-D maps, offering an efficient solution for real-time dense SLAM. More recent works like NICE-SLAM (Zhu et al., 2022) and NICER-SLAM (Zhu et al., 2024) introduced neural implicit scene encoding to achieve scalable, memory-efficient dense reconstructions. Real-time large-scale dense RGB-D SLAM (Whelan et al., 2015) further improved volumetric fusion, enabling robust reconstruction in complex and dynamic environments. ESLAM (Johari et al., 2023) leveraged photometric consistency for optimization, enhancing the accuracy of dense reconstructions. Recent innovations like Point-SLAM (Sandström et al., 2023) have integrated real-time dense mapping for large-scale scenarios, while Co-SLAM (Wang et al., 2023) jointly optimizes coordinate and sparse parametric encodings to improve reconstruction fidelity. Vox-Fusion (Yang et al., 2022) has pushed the boundaries further by incorporating voxel-based feature maps, providing efficient and detailed scene representations. DROID-SLAM (Teed and Deng, 2021) stands out for its use of recurrent neural networks (RNNs) to achieve end-to-end optimization, which makes it particularly robust in estimating both camera pose and depth maps. Despite its advancements in camera tracking accuracy and its ability to provide relatively dense point clouds, DROID-SLAM still falls short in enhancing the visual fidelity of the reconstructed maps. Its mapping output often lacks the level of detail and consistency required for high-quality scene representation, leaving room for further improvement in dense 3D reconstruction.
@@ 2.2. Differentiable Rendering
Kerbl et al. addressed this by introducing a 3D scene representation using 3D Gaussians and a fast tile-based rasterizer, achieving real-time rendering with high visual quality (3DGS) (Kerbl et al., 2023). Building on the efficiency of 3D Gaussian Splatting, DreamGaussian (Tang et al., 2023) applied it to generative 3D modeling with mesh extraction and texture refinement for text-to-3D tasks. Wu et al. further extended its use for dynamic scenes, combining 3D Gaussians with 4D neural voxels, enabling efficient feature encoding and Gaussian deformation prediction over time (Wu et al., 2024).
@@ 2.3. Differentiable Rendering SLAM
Recently, 3DGS has shown great promise in 3D reconstruction by avoiding the aforementioned limitations of NeRF-based representations. It also achieves faster rendering speeds, which is why we chose it as our primary map representation. Monocular Photo-SLAM (Huang et al., 2024) pioneered the integration of photometric consistency into SLAM frameworks, achieving real-time photorealistic mapping from monocular input. GLORIE-SLAM (Zhang et al., 2024) advances RGB-only dense SLAM with a neural point cloud representation and a novel DSPO layer for bundle adjustment, jointly optimizing pose, depth, and scale to address the lack of geometric priors in monocular SLAM. This approach improves mapping and rendering accuracy while remaining computationally efficient. RGB-D Methods like SplaTAM (Keetha et al., 2024) and GS-SLAM (Yan et al., 2024b) expand on this by incorporating both depth and color information, using 3DGS for efficient 3D scene representation. These systems enable tasks such as object interaction and robot manipulation, enhancing SLAM applications in robotics.
@@ 3.3.4. SLAM-Informed Adaptive Densification
While traditional 3D Gaussian Splatting (3DGS) densification methods excel at densifying sparse point clouds (e.g., Photo-SLAM), they are inefficient for our system, which generates dense point clouds for each frame. Such methods fail to address the early-stage inaccuracies in SLAM-generated point clouds and add unnecessary computational overhead for densification. To overcome these challenges, we propose SLAM-Informed Adaptive Densification (SIAD), which leverages SLAM’s dynamic updates to refine the point cloud online. By pruning erroneous points, incorporating new ones, and adapting Gaussian representations based on reliable masks, our approach ensures accurate and efficient point cloud refinement without redundant densification steps.
@@ 3.3.5. Keyframe Selection and Optimization Strategy
We maintain a frontend keyframe window with a default size of 25. At each iteration, the system updates the camera poses and depth estimates for all keyframes within this window using a bundle adjustment-like optimization process. These updates are subsequently propagated to the mapping module, where the dense point cloud is refined and reconstructed into a Gaussian-based scene representation.
@@
We adopt same rending method with 3DGS, in which each Gaussian is projected onto the image plane as a 2D elliptical footprint based on its mean $\mu$ and covariance matrix $\Sigma$. The rendering pipeline accumulates contributions from all Gaussians through a rasterization process, where the opacity and color of each pixel are computed by blending overlapping Gaussian footprints. This method ensures efficient rendering while preserving high visual fidelity and geometric accuracy, making it particularly suitable for real-time applications.
@@
To achieve high-quality rendering that balances photometric consistency and geometric accuracy, we define a composite loss function. The RGB loss $L_{\text{rgb}}$, inspired by Plenoxels (Yu et al., 2021) and 3DGS, ensures consistency between the rendered and ground truth images. Unlike 3DGS, which uses a combination of $L_{1}$ loss and SSIM loss, we enhance perceptual quality by incorporating multi-scale SSIM (MS-SSIM) (Wang et al., 2003), which captures perceptual consistency across different resolutions. The RGB loss is defined as:
@@
Here $L_{1}$ minimizes pixel-wise differences, and $L_{\text{ms-ssim}}$ improves perceptual quality across multiple scales. Specifically, MS-SSIM computes the structural similarity at multiple resolution levels, progressively downscaling the input images. This formulation ensures robustness to local variations and better perceptual alignment compared to the single-scale SSIM used in 3DGS.
@@ 4.1. Implementation and Experiment Setup
We evaluated SplatMap against several state-of-the-art (SOTA) SLAM systems, focusing on both RGB-D and monocular setups. For RGB-D systems, we compared with the neural radiance field (NeRF)-based SLAM systems NICE-SLAM(Zhu et al., 2022) and ESLAM, as well as Gaussian-based SLAM methods including GS-SLAM(Yan et al., 2024b) and SplaTAM(Keetha et al., 2024). For monocular systems, we selected GLORIE-SLAM(Zhang et al., 2024) and Photo-SLAM (Huang et al., 2024).
@@ 4.2. Tracking Results
Photo-SLAM (Huang et al., 2024) 1.09 0.08 1.18 0.23 0.58 0.32 0.69 0.72 0.72
@@
Table 1. Average Trajectory Error (ATE) for Different SLAM Methods.
@@ 4.3. Mapping Results
PhotoSLAM (Huang et al., 2024) PSNR 33.30 29.77 31.30 33.18 36.99 37.59 31.79 31.62 34.17
@@ 4.3.1. Replica
RGBD SplaTAM (Keetha et al., 2024) PSNR - 22.00 - - 24.50 21.90
@@
MONO MonoGS (Matsuki et al., 2024) PSNR 18.81 19.67 19.16 18.41 16.17 20.63
@@
GIORIE-SLAM (Zhang et al., 2024) PSNR 20.99 20.26 19.09 18.78 25.62 21.21
@@ 4.4. Ablation Study
Base SMOOTH SIAD MSSSIM PSNR SSIM LPIPS
$\checkmark$ $\times$ $\times$ $\times$ 33.344 0.979 0.054
$\checkmark$ $\times$ $\checkmark$ $\times$ 33.943 0.982 0.046
$\checkmark$ $\checkmark$ $\times$ $\times$ 34.444 0.985 0.037
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 34.549 0.985 0.036
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 34.693 0.986 0.033
Table 4. Ablation Study on Replica Dataset Room0
Table 4 presents the results of the ablation study on the Replica Dataset Room0, evaluating the impact of different components added to the base model. The Base model, which is an online Gaussian SLAM system with edge-aware loss, achieves a PSNR of 33.344, SSIM of 0.979, and LPIPS of 0.054. By incorporating SIAD, the PSNR improves to 33.943, and both SSIM (0.982) and LPIPS (0.046) metrics indicate enhanced image quality, demonstrating the effectiveness of SIAD in refining the reconstruction.
Adding the SMOOTH component improves the PSNR to 34.444, SSIM to 0.985, and LPIPS to 0.037, indicating that smoothness constraints significantly enhance geometric consistency and rendering quality. Combining SMOOTH and SIAD further boosts the PSNR to 34.549, while maintaining a high SSIM of 0.985 and a slightly improved LPIPS of 0.036. Finally, integrating MSSSIM (Multi-Scale Structural Similarity) yields the best results, with a PSNR of 34.693, SSIM of 0.986, and LPIPS of 0.033.
5. Conclusion

### v3

@@ Abstract.
Achieving high-fidelity 3D reconstruction from monocular video remains challenging due to the inherent limitations of traditional methods like Structure-from-Motion (SfM) and monocular SLAM in accurately capturing scene details. While differentiable rendering techniques such as Neural Radiance Fields (NeRF) address some of these challenges, their high computational costs make them unsuitable for real-time applications. Additionally, existing 3D Gaussian Splatting (3DGS) methods often focus on photometric consistency, neglecting geometric accuracy and failing to exploit SLAM’s dynamic depth and pose updates for scene refinement. We propose a framework integrating dense SLAM with 3DGS for near real-time, high-fidelity dense reconstruction. Our approach introduces SLAM-Informed Adaptive Densification, which dynamically updates and densifies the Gaussian model by leveraging dense point clouds from SLAM. Additionally, we incorporate Geometry-Guided Optimization, which combines edge-aware geometric constraints and photometric consistency to jointly optimize appearance and geometry of the 3DGS scene representation, enabling detailed and accurate SLAM mapping reconstruction. Experiments on the Replica and TUM-RGBD datasets demonstrate the effectiveness of our approach, achieving state-of-the-art results among monocular systems. Specifically, our method achieves a PSNR of 36.864, SSIM of 0.985, and LPIPS of 0.040 on Replica, representing improvements of 10.7%, 6.4%, and 49.4%, respectively, over the previous SOTA. On TUM-RGBD, our method outperforms the closest baseline by 10.2%, 6.6%, and 34.7% in the same metrics. These results highlight the potential of our framework in bridging the gap between photometric and geometric dense 3D scene representations, paving the way for practical and efficient monocular dense reconstruction. A demonstration of the results can be found in the accompanying video: https://youtu.be/Pr_kyWQQkGo.
@@
copyright: acmcopyrightjournal: PACMCGITjournalyear: 2025journalvolume: 8journalnumber: 1article: 20publicationmonth: 5doi: 10.1145/3728310ccs: Computing methodologies Reconstruction
@@ 1. Introduction
Figure 1. Visual comparison of reconstructed scenes. Our method (SplatMap) achieves superior rendering fidelity and geometric accuracy, closely matching the ground truth, while outperforming GIORIE-SLAM (Zhang et al., 2024) and Photo-SLAM (Huang et al., 2024a) in fine structural details.
@@
In order to achieve real-time 3D reconstruction with high-fidelity rendering, researchers have explored the potential of integrating SLAM with differentiable rendering (Zhu et al., 2022, 2024; Yang et al., 2022; Keetha et al., 2024; Johari et al., 2023; Sandström et al., 2023; Yan et al., 2024a; Matsuki et al., 2024; Zhang et al., 2024; Huang et al., 2024a; Peng et al., 2024). However, monocular SLAM systems, particularly during early mapping stages, are prone to producing inaccurate point clouds due to limited observations, shallow triangulation baselines, and weak pose constraints. These inaccuracies propagate into downstream mapping tasks, leading to artifacts such as ”ghosting” on walls and furniture, especially when observed from untrained viewpoints. Traditional 3DGS densification pipelines typically rely on RGB losses to correct these errors, introducing additional computational overhead as they compensate for the initial inaccuracies.
@@ 2.1. Simultaneous Localization and Mapping (SLAM)
In contrast, dense SLAM methods aim to reconstruct detailed 3D scene geometry and appearance simultaneously, often leveraging depth maps or pixel-level photometric consistency for optimization. ElasticFusion (Whelan et al., 2015) circumvented traditional pose graph optimization by directly refining dense RGB-D maps, offering an efficient solution for real-time dense SLAM. More recent works like NICE-SLAM (Zhu et al., 2022) and NICER-SLAM (Zhu et al., 2024) introduced neural implicit scene encoding to achieve scalable, memory-efficient dense reconstructions. Real-time large-scale dense RGB-D SLAM (Whelan et al., 2015) further improved volumetric fusion, enabling robust reconstruction in complex and dynamic environments. ESLAM (Johari et al., 2023) leveraged photometric consistency for optimization, enhancing the accuracy of dense reconstructions. Recent innovations like Point-SLAM (Sandström et al., 2023) have integrated real-time dense mapping for large-scale scenarios, while Co-SLAM (Wang et al., 2023) jointly optimizes coordinate and sparse parametric encodings to improve reconstruction fidelity. Vox-Fusion (Yang et al., 2022) has pushed the boundaries further by incorporating voxel-based feature maps, providing efficient and detailed scene representations. DROID-SLAM (Teed and Deng, 2021) stands out for its use of recurrent neural networks (RNNs) to achieve end-to-end optimization, which makes it particularly robust in estimating both camera pose and depth maps. Despite its advancements in camera tracking accuracy and its ability to provide relatively dense point clouds, DROID-SLAM still falls short in enhancing the visual fidelity of the reconstructed maps. Its mapping output often lacks the level of detail and consistency required for high-quality scene representation, leaving room for further improvement in dense 3D reconstruction.
@@ 2.2. Differentiable Rendering
Kerbl et al. addressed this by introducing a 3D scene representation using 3D Gaussians and a fast tile-based rasterizer, achieving real-time rendering with high visual quality (3DGS) (Kerbl et al., 2023). In addition to the aforementioned works, recent efforts by Huang et al. have demonstrated that 2D Gaussian Splatting can markedly improve the geometric accuracy of radiance field reconstruction (Huang et al., 2024b). Their approach refines the rendering process by focusing on geometric fidelity, thereby achieving more accurate and visually consistent outputs. Building on the efficiency of 3D Gaussian Splatting, DreamGaussian (Tang et al., 2023) applied it to generative 3D modeling with mesh extraction and texture refinement for text-to-3D tasks. Wu et al. further extended its use for dynamic scenes, combining 3D Gaussians with 4D neural voxels, enabling efficient feature encoding and Gaussian deformation prediction over time (Wu et al., 2024).
@@ 2.3. Differentiable Rendering SLAM
Recently, 3DGS has shown great promise in 3D reconstruction by avoiding the aforementioned limitations of NeRF-based representations. It also achieves faster rendering speeds, which is why we chose it as our primary map representation. Monocular Photo-SLAM (Huang et al., 2024a) pioneered the integration of photometric consistency into SLAM frameworks, achieving real-time photorealistic mapping from monocular input. GLORIE-SLAM (Zhang et al., 2024) advances RGB-only dense SLAM with a neural point cloud representation and a novel DSPO layer for bundle adjustment, jointly optimizing pose, depth, and scale to address the lack of geometric priors in monocular SLAM. This approach improves mapping and rendering accuracy while remaining computationally efficient. For RGB-D input systems, advancements in real-time 3D reconstruction have also been propelled by novel integration strategies of Gaussian splatting techniques. In this context, RTG-slam by Peng et al. has shown promising results by incorporating 3D Gaussian splatting into RGB-D SLAM frameworks (Peng et al., 2024). Their work not only enables efficient large-scale reconstruction but also delivers competitive performance in real-time settings, thereby providing a robust alternative to conventional methods. Other works like SplaTAM (Keetha et al., 2024) and GS-SLAM (Yan et al., 2024b) also expand on this by incorporating both depth and color information, using 3DGS for efficient 3D scene representation. These systems enable tasks such as object interaction and robot manipulation, enhancing SLAM applications in robotics.
@@ 3. Methodology
It is important to note that our approach builds upon existing DROID-SLAM (Teed and Deng, 2021) and 3DGS frameworks. For the sake of completeness, Sec.3.2 reviews the DROID-SLAM–based tracking module, and Sec.3.3.2 to  3.3.3 describe the 3DGS components inherited from prior work. In contrast, our novel contributions—including the SLAM-Informed Adaptive Densification and Geometry-Guided Optimization are detailed in Sec.3.3.4 and  3.3.5.
@@ 3.3.4. SLAM-Informed Adaptive Densification
While traditional 3DGS densification methods excel at densifying sparse point clouds (e.g., Photo-SLAM), they are inefficient for our system, which generates dense point clouds for each frame. Such methods fail to address the early-stage inaccuracies in SLAM-generated point clouds and add unnecessary computational overhead for densification. To overcome these challenges, we propose SLAM-Informed Adaptive Densification (SIAD), which leverages SLAM’s dynamic updates to refine the point cloud online. By pruning erroneous points, incorporating new ones, and adapting Gaussian representations based on reliable masks, our approach ensures accurate and efficient point cloud refinement without redundant densification steps. Our method circumvents the need for traditional cloning and splitting operations typically used in 3DGS densification, thereby streamlining the process and avoiding the associated computational overhead.
@@ 3.3.5. Keyframe Selection and Optimization Strategy
We maintain a frontend keyframe window with a default size of 25. This window size is chosen empirically to balance the need for capturing sufficient geometric diversity and controlling computational overhead. Our preliminary experiments indicated that a window size of 25 is effective in maintaining a robust optimization process while limiting the complexity of the bundle adjustment-like updates. Similar strategies have been adopted in related work, where keyframe retention is controlled via a maximum age parameter to ensure a balance between temporal context and near real-time performance. At each iteration, the system updates the camera poses and depth estimates for all keyframes within this window using a bundle adjustment-like optimization process. These updates are subsequently propagated to the mapping module, where the dense point cloud is refined and reconstructed into a Gaussian-based scene representation.
@@
We adopt the same rendering method with 3DGS, in which each Gaussian is projected onto the image plane as a 2D elliptical footprint based on its mean $\mu$ and covariance matrix $\Sigma$. The rendering pipeline accumulates contributions from all Gaussians through a rasterization process, where the opacity and color of each pixel are computed by blending overlapping Gaussian footprints. This method ensures efficient rendering while preserving high visual fidelity and geometric accuracy, making it particularly suitable for online applications.
@@
To achieve high-quality rendering that balances photometric consistency and geometric accuracy, we define a composite loss function. The RGB loss $L_{\text{rgb}}$, inspired by Plenoxels (Yu et al., 2021) and 3DGS, ensures consistency between the rendered and ground truth images. Unlike 3DGS, which uses a combination of $L_{1}$ loss and SSIM loss, we enhance perceptual quality by incorporating multi-scale SSIM (MS-SSIM) (Wang et al., 2003), which captures perceptual consistency across different resolutions. Specifically, MS-SSIM is computed by first calculating the SSIM between the rendered image and the ground truth using a fixed window size of 11. Then, the input images are progressively downsampled via 2D average pooling with a stride of 2, and the SSIM is re-evaluated at each of the three scales. The final MS-SSIM loss is obtained by averaging the SSIM values across these scales. The RGB loss is defined as:
@@
Here $L_{1}$ minimizes pixel-wise differences, and $L_{\text{ms-ssim}}$ improves perceptual quality across multiple scales. This multi-scale approach ensures robustness to local variations and yields better perceptual alignment compared to the single-scale SSIM used in 3DGS.
@@ 4.1. Implementation and Experiment Setup
We evaluated SplatMap against several state-of-the-art (SOTA) SLAM systems, focusing on both RGB-D and monocular setups. For RGB-D systems, we compared with the neural radiance field (NeRF)-based SLAM systems NICE-SLAM(Zhu et al., 2022) and ESLAM, as well as Gaussian-based SLAM methods including GS-SLAM(Yan et al., 2024b) and SplaTAM(Keetha et al., 2024). For monocular systems, we selected GLORIE-SLAM(Zhang et al., 2024) and Photo-SLAM (Huang et al., 2024a).
@@ 4.2. Tracking Results
Table 1. Average Trajectory Error (ATE) for Different SLAM Methods.
@@
Photo-SLAM (Huang et al., 2024a) 1.09 0.08 1.18 0.23 0.58 0.32 0.69 0.72 0.72
RTG-slam (Peng et al., 2024) 0.18 0.20 0.18 0.13 0.22 0.12 0.22 0.20 0.19
@@ 4.3. Mapping Results
PhotoSLAM (Huang et al., 2024a) PSNR 33.30 29.77 31.30 33.18 36.99 37.59 31.79 31.62 34.17
@@ 4.3.1. Replica
RGBD SplaTAM PSNR - 22.00 - - 24.50 21.90
@@
MONO MonoGS PSNR 18.81 19.67 19.16 18.41 16.17 20.63
@@
GIORIE-SLAM PSNR 20.99 20.26 19.09 18.78 25.62 21.21
@@ 4.3.2. TUM
4.4. Runtime Analysis
We assess the near real-time performance of our algorithm in Tab. 4. We benchmark the runtime on an NVIDIA RTX A6000 GPU with 48 GB of memory. Our system operates at an average frame rate of 3.74 fps, which, while lower than that of GO-SLAM (Zhang et al., 2023)and ESLAM (Johari et al., 2023), significantly outperforms Nice-SLAM (Zhu et al., 2022), GlORIE-SLAM (Zhang et al., 2024), and MonoGS (Matsuki et al., 2024). The peak GPU memory consumption of our method is 14.96 GiB, which is competitive with existing state-of-the-art approaches. These results highlight that our method offers a balanced trade-off between computational efficiency and mapping accuracy, making it viable for online applications.
Table 4. Memory and Running Time Evaluation on Replica room0
ESLAM Nice-SLAM GO-SLAM GlORIE-SLAM MonoGS Ours
Avg. FPS 5.55 0.48 8.36 0.23 0.32 3.74
GPU Usage (GiB) - - 18.50 15.22 14.62 14.96
@@ 4.5. Ablation Study
Table 5. Ablation Study on Replica and TUM-RGBD Datasets
Ablation Components Replica-r0 TUM-RGBD-f1/desk2
Base SMOOTH SIAD SSIM PSNR SSIM LPIPS PSNR SSIM LPIPS
A $\checkmark$ $\times$ $\times$ $\times$ 33.344 0.979 0.054 23.339 0.871 0.219
B $\checkmark$ $\times$ $\checkmark$ $\times$ 33.943 0.982 0.046 23.362 0.866 0.213
C $\checkmark$ $\checkmark$ $\times$ $\times$ 34.444 0.985 0.037 23.951 0.883 0.190
D $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 34.549 0.985 0.036 24.104 0.888 0.191
E $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 34.693 0.986 0.033 24.413 0.896 0.179
Figure 6. Ablation Study Visualization. The labels correspond to the respective experimental settings in Tab. 5
Table 5 presents the results of the ablation study on the Replica Dataset Room0 and TUM-RGBD dataset (f1/desk2), evaluating the impact of different components added to the base model. The Base model in our work refers to a vanilla online Gaussian SLAM system that integrates DROID-SLAM with 3DGS and incorporates an edge-aware loss to improve structural consistency. Here, “SMOOTH” refers to our proposed smooth weighting function (introduced in Sec. 3.3.5, Equ. 13), which modulates the edge-aware loss using a Gaussian-like function to balance edge emphasis with spatial continuity. The configuration denoted as ”Base + SIAD + SMOOTH + MS-SSIM” represents the full experimental setup used throughout our paper. In particular, this exact configuration is employed to produce the results shown in Fig. 4 and Fig. 5
The Base model achieves a PSNR of 33.344, SSIM of 0.979, and LPIPS of 0.054. By incorporating SIAD, the PSNR improves to 33.943, and both SSIM (0.982) and LPIPS (0.046) metrics indicate enhanced image quality, demonstrating the effectiveness of SIAD in refining the reconstruction. Adding the SMOOTH component improves the PSNR to 34.444, SSIM to 0.985, and LPIPS to 0.037, indicating that smoothness constraints significantly enhance geometric consistency and rendering quality. Combining SMOOTH and SIAD further boosts the PSNR to 34.549, while maintaining a high SSIM of 0.985 and a slightly improved LPIPS of 0.036. Finally, integrating MS-SSIM (Multi-Scale Structural Similarity) yields the best results, with a PSNR of 34.693, SSIM of 0.986, and LPIPS of 0.033.
The experiments on the TUM-RGBD dataset exhibit a similar trend: incorporating SIAD, SMOOTH, and MS-SSIM consistently improves the reconstruction quality. This consistency across two distinct datasets demonstrates that the improvements observed in our ablation study are generalizable and effective across diverse scenes.
The observed improvements in the ablation study are a direct consequence of our architectural innovations. Specifically, the SIAD module refines the dense point cloud by leveraging SLAM’s dynamic updates to prune erroneous points and densify only those that are newly reliable, thereby reducing ghosting artifacts and improving geometric accuracy. The SMOOTH component—our proposed smooth weighting function—modulates the edge-aware loss to prevent over-penalization in high-gradient regions while maintaining spatial continuity, which enhances the overall rendering consistency. Moreover, the incorporation of MS-SSIM into the composite RGB loss ensures multi-scale perceptual alignment, capturing structural details more robustly than single-scale SSIM. Together, these components synergistically contribute to significant gains in PSNR, SSIM, and LPIPS, as demonstrated consistently across both the Replica and TUM-RGBD datasets.
5. Conclusion and Future Work